OpenStudy (anonymous):

Write the expression as one logarithm. 2log (y^8/x) - 6log y + 1/2 log (x^4 y^2)

OpenStudy (anonymous):

you need to know two properties of logarithms. Ready? \[\log_{n}ab = \log_{n}a+\log_{n}b \] \[\log_{n}a ^{x} = x \log_na\]

OpenStudy (anonymous):

\[2\log(\frac{y^8}{x})-6\log(y)+\frac{1}{2} \log(x^4 y^2)\]

OpenStudy (anonymous):

if that is the question there is a whole raft of cancellation and i think you are just left with \[11 \log(y)\]

OpenStudy (anonymous):

wanna do it step by step?